the speed of a paricle moving in a cirle of radius 61/2  m is increasing 3 m/s2 . find net acceleration at t=2 sec ( Given, v=0, at t=0 )

Grade:12th pass

1 Answers

Khimraj
5 years ago
tangential acceleration = 3 m/s2
velocity after 2 sec = 3*2 = 6 m/s
so radial acceleration = v2/r =6\sqrt{6} m/s2
ao total acceleration = \sqrt{3^{2} + (6\sqrt{6})^{2}} = 15 m/s^2
Hopeit clears If yo like answer then please approve it.